Two wine merchants enter Paris, one of them with 64 casks of wine, the other with 20. Since they have not enough money to pay the custom duties, the first pays 5 casks of wine and 40 francs, and the second pays 2 casks of wine and receives 40 francs in change. What is the price of each cask of wine and the duty on it?
Answer by richwmiller(17219)

64*d=5c+40f
20*d=2c-40f
c=120f d=10f
check 20
20*120=2400
200f
240-200=40
check 64
5*120=600
64*10=640
640-600=40